If y = 2 2 ^x , then show that dy/dx = y( log 2 ) 2 *2 x

If y = 22^x, then show that dy/dx = y( log 2 )*2x 
 
 

Grade:11

1 Answers

kkbisht
5 years ago
Let y=22^x 
Take log  we obtain
logy = 2x log2 
on differetiation wrt x we get
(1/y) dy/dx= log 2 .( 2x Log2)= ( log2)2 2x   = > dy/dx= y (log)2 .2x
 
kkbisht